# Evaluating multiple models using labeled and unlabeled data

### v2

@@ D.3 Comparison to baselines drawn from weak supervision
Popular approaches to weak supervision including Snorkel (Ratner et al., 2017) and FlyingSquid (Fu et al., 2020) implement a latent variable model equivalent to Dawid-Skene. Both works build on Dawid-Skene to incorporate information about pairwise correlations between labeling functions; (Ratner et al., 2017) employs a technique to infer dependencies, while (Fu et al., 2020) assume these dependencies to be user-provided. When we applied a standard approach to dependency inference (Bach et al., 2017) in our setting, we observed that (1) all classifiers are inferred to be dependent on one another, and (2) the number of dependencies raised issues with convergence. It is thus not feasible to incorporate dependency inference, and the resulting latent variable model is equivalent to Dawid-Skene.
D.4 Comparison to ensembling
@@
While we limit the scope of our experiments in the main text to semi-supervised methods that make use of both labeled and unlabeled data, another approach would be to produce an estimate of $Pr(y=k|s^{(i)})$ by averaging the classifier scores. This approach results in an unbiased metric estimator when the resulting ensemble is calibrated, as theoretical results by Ji et al. (2020) show. Such an approach has natural downsides: it is sensitive to the composition of the classifier set, does not improve with the introduction of labeled data, and relies on an assumption of ensemble calibration that is unlikely to hold in practice (Wu & Gales, 2021). Here, we provide experiments to illustrate this behavior.
@@
Using the semisynthetic setting described in Section 6.4, we artificially increase the expected calibration error of each classifier using a generalized logistic function parameterized by $a$. Specifically, we transform classifier score $s$ to be $\frac{s^{a}}{s^{a}+(1-s)^{a}}$, effectively increasing overconfidence for higher $s$ and increasing underconfidence for lower $s$. As in the semisynthetic experiments, we generate 500 semisynthetic classifier sets, where each classifier in a set is trained on 100 examples distinct from the training data for other classifiers in the set (results are robust to this choice of training dataset size). Each set contains three classifiers.
@@
Figure S3 reports our results. As the average calibration among classifiers in a set varies, SSME consistently improves over the use of an ensemble. This aligns with our intuition, and indicates the value of using labeled data in conjunction with unlabeled data. Interestingly, miscalibration has little effect on the ensemble when estimating AUPRC; here, SSME and ensembling perform similarly.
@@
D.5 Discussion of classifier correlation
[…]
@@ Appendix E Method Details
E.1 Metric Estimation
Given a vector $p\in\Delta^{K-1}$ over $K$ classes, let $\mathbf{s}=\text{ALR}(\mathbf{p})=\left[\log\frac{\mathbf{p}_{1}}{\mathbf{p}_{K}},\log\frac{\mathbf{p}_{2}}{\mathbf{p}_{K}},\cdots,\log\frac{\mathbf{p}_{K-1}}{\mathbf{p}_{K}}\right]\in\mathbb{R}^{K-1}$. To invert, $\mathbf{p}_{i}=\frac{e^{\mathbf{s}_{i}}}{1+\sum_{k=1}^{K-1}e^{\mathbf{s}_{k}}}$ for $i<K$ and $\mathbf{p}_{K}=\frac{1}{1+\sum_{k=1}^{K-1}e^{\mathbf{s}_{k}}}$. The ALR transform maps unit-sum data into real space, where it is easier to fit mixture models. The inverse allows us to map samples from the mixture model in real space back to the simplex $\Delta^{K-1}$. For details, see Pawlowsky-Glahn & Buccianti (2011).
@@ E.2 Metric Estimation
SSME is able to estimate any metric that is a function of the classifier probabilities $p$ and label $y$. We approximate the joint distribution $P(y,\mathbf{p})$ with a mixture model model $P_{\theta}(y,\mathbf{s})$, where $\mathbf{s}$ refers to the ALR-transformed classifier probabilities (i.e. “classifier scores”) [^2] Recall that ALR is a bijection, so we use the inverse mapping $\text{ALR}^{-1}:\mathbb{R}^{K-1}\rightarrow\Delta^{K-1}$ to transform our mixture distribution in real space back to probability space.. We refer to $P(y,\mathbf{p})$ for ease of notation in this section; it is equivalent, through invertible mapping, to $P(y,\mathbf{s})$.
@@
In other words, we compute the expectation that the true label agrees with the predicted label for each point . This expectation is $\mathbf{p}^{(i)}$. This expectation is computed over $P_{\theta}(y|\mathbf{p})$ One can interpret $P_{\theta}(y|\mathbf{p})$ as a “recalibration” step: given a set of classifier guesses $\mathbf{p}$, what is the true distribution of $y$?
@@
Expected Calibration Error (ECE) measures the alignment between a model’s predicted probabilities $\mathbf{p}_{j}$ and the ground truth labels $y$. In particular, ECE compares the model’s reported confidence to the true class likelihoods, averaged over the dataset. We write out our ECE estimation procedure for the binary case, and it extends readily to definitions of calibration in multiclass settings (Gupta & Ramdas, 2022). Binary ECE can be written as:
@@
Then, to approximate the ECE with the datasets $\{\mathbf{p}^{i}\}_{i=1}^{n_{u}}$ and $\{\mathbf{p}^{i},y^{(i)}\}_{i=1}^{n_{\ell}}$, one can sample $y^{(i)}\sim P_{\theta}(y|\mathbf{p}^{(i)})$ for each unlabeled sample $i$ and then use the standard histogram binning procedure (Guo et al., 2017) using both the true labels for the labeled dataset and the sampled labels for the unlabeled dataset. In this approach, we treat the sampled labels $y^{(i)}$ as true labels for unlabeled examples. To ensure our procedure is robust against sampling noise, we draw samples of $y^{(i)}$ repeatedly for a fixed number of draws (500). We then compute ECE separately for each of these 500 draws and average ECE across all draws.
Alternatively, one could also directly use $P_{\theta}(y|\mathbf{p})$ to estimate ECE. In particular, we can write:
$\text{ECE}_{j}\approx\frac{1}{n_{u}+n_{\ell}}\sum_{i=1}^{n_{u}+n_{\ell}}\left|P_{\theta}\left(y=1|\mathbf{p}^{(i)}_{j}\right)-\mathbf{p}^{(i)}_{j}\right|$
In this approach, we don’t sample the labels $y$ for unlabeled examples but instead directly use $P_{\theta}(y|\mathbf{p})$, which provides us (an estimate of) the true distribution of $y$. Instead, we directly use our estimate for the conditional label distribution $P_{\theta}\left(y=1|\mathbf{p}^{(i)}_{j}\right)$. In our experiments, we use the first approach described, i.e. sampling $y^{(i)}$ for unlabeled examples and then using the standard binning and averaging procedure.
@@
E.3 EM Algorithm
[…]
We use the EM algorithm to fit SSME, which iterates between the $E$-step and $M$-step updates, described below.
E-step: For the $t$th update, we compute:
$P^{t+1}_{\theta}(y^{(i)}=k|\mathbf{s}^{(i)})=\frac{p^{t}(y=k)P_{\theta}^{t}(\mathbf{s}^{(i)}|y^{(i)}=k)}{\sum_{\ell=1}^{K}p^{t}(y=\ell)P_{\theta}^{t}(\mathbf{s}^{(i)}|y^{(i)}=\ell)}$
This is combined with the prior $p^{t}(y=k)$ to produce posteriors $P^{t+1}_{\theta}(y^{(i)=k|\mathbf{s}^{(i)})}$. We then fix the labels for labeled example, setting $P_{\theta}(P_{\theta}(y^{(i)=k^{*}|\mathbf{s}^{(i)})}=1$ for the correct class $k*$ and to 0 for all other classes.
M-step: During the M-step, we update the prior $p^{t}(y=k)$. To do so, we calculate:
$p^{t+1}(y=k)=\frac{1}{n_{u}+n_{\ell}}\left(\sum_{i=1}^{n_{\ell}}I(y^{(i)}=k)+\sum_{i=1}^{n_{u}}P^{t+1}_{\theta}(y^{(i)}=k|\mathbf{s}^{(i)})\right)$
[…]
We alternate between the above two updates for 1000 epochs.

### v3

@@ D.9 Comparison to baselines drawn from weak supervision
Popular approaches to weak supervision including Snorkel [58] and FlyingSquid [27] implement a latent variable model equivalent to Dawid-Skene. Both works build on Dawid-Skene to incorporate information about pairwise correlations between labeling functions; [58] employs a technique to infer dependencies, while [27] assume these dependencies to be user-provided. When we applied a standard approach to dependency inference [4] in our setting, we observed that (1) all classifiers are inferred to be dependent on one another, and (2) the number of dependencies raised issues with convergence. It is thus not feasible to incorporate dependency inference, and the resulting latent variable model is equivalent to Dawid-Skene.
D.10 Comparison to ensembling
@@
While we limit the scope of our experiments in the main text to semi-supervised methods that make use of both labeled and unlabeled data, another approach would be to produce an estimate of $Pr(y=k|s^{(i)})$ by averaging the classifier scores. This approach results in an unbiased metric estimator when the resulting ensemble is calibrated, as theoretical results by [37] show. Such an approach has natural downsides: it is sensitive to the composition of the classifier set, does not improve with the introduction of labeled data, and relies on an assumption of ensemble calibration that is unlikely to hold in practice [71]. Here, we provide experiments to illustrate this behavior.
We use a semisythetic setting to conduct a comparison to ensembling. To do so, we create sets of three classifiers based on the widely-used Adult dataset [7], where the task is to predict whether a person’s income is above \$50K. To create differences between the three classifiers in a set, we train them on random fixed-size samples of 50 labeled examples from different portions of the dataset, partitioned based on age. In doing so, our semi-synthetic classifier sets mimic how training data for different real-world classifiers can differ in meaningful ways. We repeat this procedure to produce 500 sets of three classifiers, where sets differ in the training data provided to each classifier. Our procedure naturally produces random variation in classifier properties, like accuracy and calibration. As with the real datasets, we produce three splits: a training split to learn the classifiers (50 examples), an estimation split for the performance estimation methods (20 labeled examples and 1000 unlabeled examples), and an evaluation split to measure ground truth values for each metric (10,000 examples). Each classifier is a logistic regression with default L2 regularization.
@@
We artificially increase the expected calibration error of each classifier using a generalized logistic function parameterized by $a$. Specifically, we transform classifier score $s$ to be $\frac{s^{a}}{s^{a}+(1-s)^{a}}$, effectively increasing overconfidence for higher $s$ and increasing underconfidence for lower $s$. As in the semisynthetic experiments, we generate 500 semisynthetic classifier sets, where each classifier in a set is trained on 100 examples distinct from the training data for other classifiers in the set (results are robust to this choice of training dataset size). Each set contains three classifiers.
@@
Figure S3 reports our results. As the average calibration among classifiers in a set varies, SSME consistently improves over the use of an ensemble. This aligns with our intuition, and indicates the value of using labeled data in conjunction with unlabeled data. Miscalibration has little effect on the ensemble when estimating AUPRC; here, SSME and ensembling perform similarly.
Figure S2: SSME performance in response to additional unlabeled data, additional classifiers, and improved classifiers. As indicated by our theoretical results, SSME benefits from increased amounts of unlabeled data, number of classifiers, and classifier performance.
@@
D.11 Comparison of computational cost
Fitting SSME is computationally cheap and faster than the best baseline. For every dataset and experiment configuration reported in the paper, SSME can be fit in under 5 minutes, using only 1 CPU. SSME inherits the computational complexity of kernel density estimators ($O(n^{2})$) and EM. In a matched comparison (using 20 labeled examples, 1000 unlabeled examples, and 7 classifiers) the next best-performing baseline (Bayesian-Calibration) takes on average 32.1 seconds to estimate performance, while SSME takes 21.5 seconds.
@@ Appendix E Method Details
E.1 Additive Log-Ratio Transform
Given a vector $p\in\Delta^{K-1}$ over $K$ classes, let $\mathbf{s}=\text{ALR}(\mathbf{p})=\left[\log\frac{\mathbf{p}_{1}}{\mathbf{p}_{K}},\log\frac{\mathbf{p}_{2}}{\mathbf{p}_{K}},\cdots,\log\frac{\mathbf{p}_{K-1}}{\mathbf{p}_{K}}\right]\in\mathbb{R}^{K-1}$. To invert, $\mathbf{p}_{i}=\frac{e^{\mathbf{s}_{i}}}{1+\sum_{k=1}^{K-1}e^{\mathbf{s}_{k}}}$ for $i<K$ and $\mathbf{p}_{K}=\frac{1}{1+\sum_{k=1}^{K-1}e^{\mathbf{s}_{k}}}$. The ALR transform maps unit-sum data into real space, where it is easier to fit mixture models. The inverse allows us to map samples from the mixture model in real space back to the simplex $\Delta^{K-1}$. For details, see [53].
@@ E.2 Metric Estimation
SSME is able to estimate any metric that is a function of the classifier probabilities $p$ and label $y$. We approximate the joint distribution $P(y,\mathbf{p})$ with a mixture model model $P_{\theta}(y,\mathbf{s})$, where $\mathbf{s}$ refers to the ALR-transformed classifier probabilities (i.e. “classifier scores") [^2] Recall that ALR is a bijection, so we use the inverse mapping $\text{ALR}^{-1}:\mathbb{R}^{K-1}\rightarrow\Delta^{K-1}$ to transform our mixture distribution in real space back to probability space.. We refer to $P(y,\mathbf{p})$ for ease of notation in this section; it is equivalent, through invertible mapping, to $P(y,\mathbf{s})$.
Input: Labeled set $\mathcal{D}_{\ell}=\{(\mathbf{s}^{(i)},y^{(i)})\}_{i=1}^{n_{\ell}}$; Unlabeled set $\mathcal{D}_{u}=\{\mathbf{s}^{(i)}\}_{i=1}^{n_{u}}$; $\mathbf{s}^{(i)}$ are ALR-transformed classifier scores; Fitted model $P_{\theta}(y\mid\mathbf{s})$; metric function $g(\mathbf{p},y)$; Number of samples $S$ = 500
Init: $\texttt{sum}\leftarrow 0$;
1 for $s\leftarrow 1$ to $S$ do
2 foreach $\mathbf{s}^{(i)}\in\mathcal{D}_{u}$ do
    $\tilde{y}^{(i)}\sim P_{\theta}(y\mid\mathbf{s}^{(i)})$; ;
    // Sample true label $y$ from mixture model
3
4 $\mathcal{D}^{(s)}\leftarrow\{(\mathbf{s}^{(i)},\tilde{y}^{(i)})\}\cup\mathcal{D}_{\ell}$;
5 $\texttt{metric}^{(s)}\leftarrow(n_{u}+n_{\ell})^{-1}\sum_{(\mathbf{s},y)\in\mathcal{D}^{(s)}}g(\mathrm{ALR}^{-1}(\mathbf{s}),y)$;
6 $\texttt{sum}\leftarrow\texttt{sum}+\texttt{metric}^{(s)}$;
Output: $\widehat{\texttt{metric}}=\texttt{sum}/S$ ;
// Sample-based estimate of metric
7
Algorithm 2 Sampling-based metric estimation with SSME
@@ 6 $\texttt{sum}\leftarrow\texttt{sum}+\texttt{metric}^{(s)}$;
In other words, we compute the expectation that the true label agrees with the predicted label for each point . This expectation is $\mathbf{p}^{(i)}$. This expectation is computed over $P_{\theta}(y|\mathbf{p})$ One can interpret $P_{\theta}(y|\mathbf{p})$ as a “recalibration" step: given a set of classifier guesses $\mathbf{p}$, what is the true distribution of $y$?
@@
Expected Calibration Error (ECE) measures the alignment between a model’s predicted probabilities $\mathbf{p}_{j}$ and the ground truth labels $y$. In particular, ECE compares the model’s reported confidence to the true class likelihoods, averaged over the dataset. We write out our ECE estimation procedure for the binary case, and it extends readily to definitions of calibration in multiclass settings [32]. Binary ECE can be written as:
@@
Then, to approximate the ECE with the datasets $\{\mathbf{p}^{i}\}_{i=1}^{n_{u}}$ and $\{\mathbf{p}^{i},y^{(i)}\}_{i=1}^{n_{\ell}}$, one can sample $y^{(i)}\sim P_{\theta}(y|\mathbf{p}^{(i)})$ for each unlabeled sample $i$ and then use the standard histogram binning procedure [31] using both the true labels for the labeled dataset and the sampled labels for the unlabeled dataset. In this approach, we treat the sampled labels $y^{(i)}$ as true labels for unlabeled examples. To ensure our procedure is robust against sampling noise, we draw samples of $y^{(i)}$ repeatedly for a fixed number of draws (500). We then compute ECE separately for each of these 500 draws and average ECE across all draws.
We use this first approach, but alternatively, one could also directly use $P_{\theta}(y|\mathbf{p})$ to estimate ECE. In particular, we can write:
[…]
In this approach, we don’t sample the labels $y$ for unlabeled examples but instead directly use $P_{\theta}(y|\mathbf{p})$, which provides us (an estimate of) the true distribution of $y$. Instead, we directly use our estimate for the conditional label distribution $P_{\theta}\left(y=1|\mathbf{p}^{(i)}_{j}\right)$.
In our experiments, we use the first approach described, i.e. sampling $y^{(i)}$ for unlabeled examples and then using the standard binning and averaging procedure.
@@
E.3 Alternate parameterizations
One alternative parameterization is to use a normalizing flow to model our mixture of distributions. Normalizing flows learn and apply an invertible transform $f_{\theta}$ to a random variable $\mathbf{z}\sim D_{1}$ to obtain $f_{\theta}(\mathbf{z})\sim D_{2}$. Here, we set $\mathbf{z}\sim D_{1}$ to a Gaussian mixture model and learn a transformation such that $f_{\theta}(\mathbf{z})\overset{\text{dist.}}{\approx}\mathbf{s}$, i.e., the transformed distribution roughly matches our classifier score distribution. By modeling $\mathbf{z}$ explicitly as a Gaussian mixture model, one can move back and forth between the two distributions, as $f_{\theta}^{-1}(f_{\theta}(\mathbf{z}))=\mathbf{z}$, where $f_{\theta}^{-1}$ is the inverse of $f$. Specifically, we set the distribution of $\mathbf{Z}$ to follow a Gaussian mixture:
$\mathbf{Z}|(Y=k)\sim\mathcal{N}(\mu_{k},\Sigma_{k})$
Thus, the marginal distribution of $\mathbf{Z}$ is $p_{\mathbf{Z}}(\mathbf{z})=\sum_{k=1}^{K}\mathcal{N}(\mathbf{z}|\mu_{k},\Sigma_{k})\cdot p(y=k)$ is the overall density of $\mathbf{z}$. We apply our invertible transformation $f_{\theta}$ to obtain $\mathbf{s}=f_{\theta}(\mathbf{z})$. To find $p(\mathbf{s}|y=k)$, we follow the approach of [35]:
We use the RealNVP architecture [22] to parameterize $f_{\theta}$ using 10 coupling layers, 3 fully-connected layers, and a hidden dimension of 128 between the fully connected layers. Our normalizing flow is lightweight and trains in less than a minute for each dataset in our experiments section using 1 80GB NVIDIA A100 GPU.